QUESTION 10 Find the midrange for the given sample data. The heights (in feet) of 15 different poplar trees are shown below. Find the midrange. 29.3 343 323 273 30.3 28.3 313 35.3 323 36.3 29.3 25.3 31.3 323 37.3 32.80 feet O 4.17 feet O31.3 feet 14.65 feet Click Save and Submit to save and submit. Click Save All Answers to save all answers PIC
Answer #1

Midrange = (High value + Low low)/2

Here, high value = 37.3 and low value = 25.3

Therefore, midrange of the given sample data is,

Midrange = (37.3 + 25.3)/2 = 62.6/2 = 31.3

Midrange = 31.3 feet